# Robust Score-Based Quickest Change Detection

### v1

@@ I Introduction
In contrast with the work in [17], we define the notion of least favorable distribution using Fisher divergence and provide a method to effectively identify the least favorable distribution for the post-change model.
@@
    We propose a new robust score-based quickest change detection algorithm that can be applied to score-based models and unnormalized models, namely, statistical models whose density involves an unknown normalizing constant. Specifically, we use the Hyvärinen score ([16]) to propose a robust score-based variant of the SCUSUM algorithm from [15] which we refer to as RSCUSUM. In this variant and its subsequent theory, the role of Kullback-Leibler divergence in classical change detection is replaced with the Fisher divergence between the pre-and post-change distributions. This algorithm is introduced in Section V.
@@
    From a theoretical perspective, unlike the CUSUM algorithm that leverages the fact that the likelihood ratios form a martingale under the pre-change model [9, 18], the RSCUSUM algorithm is a score-based algorithm where cumulative scores do not enjoy a standard martingale characterization. Our analysis of the delay and false alarm analysis for RSCUSUM is based on new analysis techniques. This analysis is presented in Section VI.
@@ II Problem Formulation
$\text{ARL}\overset{\Delta}{=}\mathbb{E}_{\infty}[T].$
@@
$\min_{T}\;\sup_{P_{\infty}\in\mathcal{G}_{\infty},P_{1}\in\mathcal{G}_{1}}\mathcal{L}_{\texttt{WADD}}(T)\;\quad\text{subject to}\;\quad\mathbb{E}_{\infty}[T]\geq\gamma,$ (2)
where $\gamma$ is a constraint on the ARL. The delay $\mathcal{L}_{\texttt{WADD}}$ in the above problem is a function of the true pre- and post-change laws $P_{\infty},P_{1}$ and should be designated as $\mathcal{L}_{\texttt{WADD}}^{P_{\infty},P_{1}}.$ We will, however, suppress this notation and simply refer to $\mathcal{L}_{\texttt{WADD}}^{P_{\infty},P_{1}}$ by $\mathcal{L}_{\texttt{WADD}}$. Thus, the goal in this problem is to find a stopping time $T$ to minimize the worst-case detection delay, subject to a constraint $\gamma$ on $\mathbb{E}_{\infty}[T]$. We are also interested in the minimax formulation:
$\min_{T}\;\sup_{P_{\infty}\in\mathcal{G}_{\infty},P_{1}\in\mathcal{G}_{1}}\mathcal{L}_{\texttt{CADD}}(T)\;\quad\text{subject to}\;\quad\mathbb{E}_{\infty}[T]\geq\gamma.$ (3)
@@ IV Quickest Change Detection in Unnormalized and Score-Based Models
whenever it can be well defined. We assume that the Hyvärinen scores is well defined for all $P\in\mathcal{G}_{\infty}\cup\mathcal{G}_{1}$ by Assumption V.3. Here, $\|\cdot\|_{2}$ denotes the Euclidean norm, $\nabla_{x}$ and $\Delta_{x}=\sum_{i=1}^{d}\frac{\partial^{2}}{\partial x_{i}^{2}}$ respectively denote the gradient and the Laplacian operators acting on $X=(x_{1},\cdots,x_{d})^{\top}$.
@@
$\mathbb{D}_{\texttt{F}}(P\|Q)\overset{\Delta}{=}\mathbb{E}_{X\sim P}\left[\left\|\nabla_{{x}}\log p(X)-\nabla_{{x}}\log q(X)\right\|_{2}^{2}\right],$ (8)
@@
The following theorem is established in [15] regarding the consistency and the performance of the SCUSUM algorithm.
Theorem IV.3 ([15]).
Assume that the densities $p_{\infty}$ and $p_{1}$ satisfy the regularity conditions mentioned in Assumptions II.2 and II.3. Then the following statements are true:
@@ V Robust Quickest Change Detection in Unnormalized and Score-Based Models
$Q_{\infty},Q_{1}=\operatorname*{arg\,min}_{R_{\infty}\in\mathcal{G}_{\infty},R_{1}\in\mathcal{G}_{1}}\mathbb{D}_{F}(R_{1}\|R_{\infty}).$ (14)
@@
We impose the following conditions on the classes $\mathcal{G}_{\infty}$ and $\mathcal{G}_{1}$. All lemmas and theorems in the paper, unless otherwise stated, will be assumed to be valid when all these conditions are satisfied.
Assumption V.2.
$\mathcal{G}_{\infty},\mathcal{G}_{1}$ are disjoint and are each convex.
Assumption V.3.
For each $P\in\mathcal{G}_{\infty}\cup\mathcal{G}_{1}$, mild regularity conditions introduced in [16] hold.
Assumption V.4.
@@ VI Consistency and Delay and False Alarm Analysis of the RSCUSUM Algorithm
In this section, we prove the consistency (the ability to detect the change with probability one) and provide delay and false alarm analysis of the RSCUSUM algorithm. In Section VI-A, we prove an important lemma that can be interpreted as a reverse triangle inequality for Fisher divergence. This lemma is then used in Section VI-B to prove the consistency of the RSCUSUM algorithm. In Section VI-C, we obtain a lower bound on the mean time to a false alarm, and in Section VI-D, we obtain an expression for the average detection delay of the RSCUSUM algorithm. We first make another fundamental assumption:
@@
If $\mathcal{G}_{\infty}$ contains only one element ($\mathcal{G}_{\infty}=\{P_{\infty}\}$), then clearly the least favorable pre-change distribution $Q_{\infty}$ is equal to $P_{\infty}$ and $\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})=0$. Then, Assumption VI.1 follows from the disjointness of $\mathcal{G}_{\infty},\mathcal{G}_{1}$ given in Assumption V.2 and the regularity conditions given in Assumption V.3.
@@
$\mathcal{P}_{m}=\{P_{i},\;i=1,\dots,m:\;P_{i}\in\mathcal{P}\},\;m\in\mathbb{N}^{+}$
and further suppose that $\mathcal{G}_{\infty}$ is defined to be the convex hull of this finite set:
@@ VI-A Reverse Triangle Inequality for Fisher Divergence
$=\int\big{\|}\nabla\log\bigl{(}\xi q_{1}+(1-\xi)r_{1}\bigr{)}-\nabla\log q_{\infty}\big{\|}^{2}$
@@
$\mathcal{L}^{\prime}(\xi)=\int\bigl{(}q_{1}-r_{1}\bigr{)}\big{\|}\nabla\log q_{\xi}-\nabla\log q_{\infty}\big{\|}^{2}dx+\int 2q_{\xi}\nabla\left(\frac{q_{1}-r_{1}}{q_{\xi}}\right)^{T}\bigl{(}\nabla\log q_{\xi}-\nabla\log q_{\infty}\bigr{)}dx.$
@@
$\mathcal{L}^{\prime}(1^{-})=\int\bigl{(}q_{1}-r_{1}\bigr{)}\big{\|}\nabla\log q_{1}-\nabla\log q_{\infty}\big{\|}^{2}dx\quad+\int 2q_{1}\nabla\left(\frac{q_{1}-r_{1}}{q_{1}}\right)^{T}\bigl{(}\nabla\log q_{1}-\nabla\log q_{\infty}\bigr{)}dx$
$=\mathbb{D}_{\texttt{F}}\left(Q_{1}\|Q_{\infty}\right)-\int\underbrace{r_{1}\big{\|}\nabla\log q_{1}-\nabla\log q_{\infty}\big{\|}^{2}}_{\text{term 1}}+\underbrace{2q_{1}\nabla\left(\frac{q_{1}-r_{1}}{q_{1}}\right)^{T}\bigl{(}\nabla\log q_{1}-\nabla\log q_{\infty}\bigr{)}}_{\text{term 2}}dx.$ (19)
@@
$r_{1}\big{\|}\nabla\log q_{1}-\nabla\log q_{\infty}\big{\|}^{2}$
$=r_{1}\big{\|}\nabla\log q_{1}-\nabla\log r_{1}\big{\|}^{2}+r_{1}\big{\|}\nabla\log r_{1}-\nabla\log q_{\infty}\big{\|}^{2}+\underbrace{2r_{1}\bigl{(}\nabla\log q_{1}-\nabla\log r_{1}\bigr{)}^{T}\bigl{(}\nabla\log r_{1}-\nabla\log q_{\infty}\bigr{)}}_{\text{term 1(a)}}.$ (20)
@@
$2q_{1}\nabla\left(\frac{q_{1}-r_{1}}{q_{1}}\right)^{T}\bigl{(}\nabla\log q_{1}-\nabla\log q_{\infty}\bigr{)}=2r_{1}\bigl{(}\nabla\log q_{1}-\nabla\log r_{1}\bigr{)}^{T}\bigl{(}\nabla\log q_{1}-\nabla\log q_{\infty}\bigr{)}.$ (23)
@@ VI-C False Alarm Analysis of the RSCUSUM Algorithm
We emphasize that $\rho$ is a quantity that depends upon $P_{\infty},Q_{\infty}$, and $Q_{1}$ but which does not depend upon $\omega$. For any choice of $\mathcal{G}_{\infty},\mathcal{G}_{1},P_{\infty}$ that satisfies the assumptions of this paper, we can show the existence of a $\rho$ that is a solution to 25.
@@
Theorem VI.7 implies that the ARL increases at least exponentially as the stopping threshold $\omega$ increases.
@@ VI-D Delay Analysis of the RSCUSUM Algorithm
Theorem VI.8 implies that the expected detection delay (EDD) increases at most linearly as the stop threshold $\omega$ increases for large values of $\omega$.
@@ VII Identification of the Least Favorable Distributions
$\mathbb{E}_{P_{*}}\biggl{\{}\|\nabla_{x}\log p_{*}(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{\}}=\min_{p\in\mathcal{A}_{m},\phi\in\mathcal{B}_{m}}\mathbb{E}_{P}\biggl{\{}\|\phi(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{\}}.$ (29)
@@
$\mathbb{E}_{P_{*}}\biggl{\{}\|\nabla_{x}\log p_{*}(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{\}}=\min_{p\in\mathcal{A}_{m}}\mathbb{E}_{P}\biggl{\{}\|\nabla_{x}\log p(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{\}}.$
@@
$\mathbb{E}_{P}\biggl{\{}\|\nabla_{x}\log p(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{\}}=\mathbb{E}_{P}\biggl{\{}\biggl{\|}\frac{\nabla_{x}p(X)}{p(X)}-\nabla_{x}\log r(X)\biggr{\|}_{2}^{2}\biggr{\}}$
$=\mathbb{E}_{P}\biggl{\{}\biggl{\|}\frac{\sum_{i=1}^{m}w_{i}\nabla_{x}p_{i}(X)}{\sum_{i=1}^{m}w_{i}p_{i}(X)}-\nabla_{x}\log r(X)\biggr{\|}_{2}^{2}\biggr{\}}=\mathbb{E}_{P}\biggl{\{}\biggl{\|}\sum_{i=1}^{m}u_{i}(X)\nabla_{x}\log p_{i}(X)-\nabla_{x}\log r(X)\biggr{\|}_{2}^{2}\biggr{\}},$
@@
Next, we provide a method to find the LFD in a class of Gaussian mixture models.
Theorem VII.2.
Let $\mathbb{M}_{\infty},\mathbb{M}_{1}\subset\mathbb{R}^{d}$ be disjoint, convex, and compact sets in $d-$dimensional Euclidean space. Fix a symmetric, positive-definite matrix $V\in\mathbb{R}^{d\times d}$. For $\theta\in\mathbb{M}_{\infty}\cup\mathbb{M}_{1}$, let $G_{\theta}$ be the Gaussian distribution with density $g_{\theta}$ parameterized by covariance matrix $V$ and mean $\theta:G_{\theta}=\mathcal{N}(\theta,V)$. Let the distribution class $\mathcal{G}_{\infty}$ be the convex hull of all Gaussian $\{G_{\theta}:\theta\in\mathbb{M}_{\infty}\}$ and let $\mathcal{G}_{1}$ be the corresponding convex hull of $\{G_{\theta}:\theta\in\mathbb{M}_{1}\}$. Define the $V$-norm of a vector $\theta$ as $\|\theta\|_{V}\overset{\Delta}{=}\sqrt{\theta^{T}V^{-2}\theta}$. Then, $\mathcal{G}_{\infty},\mathcal{G}_{1}$ are disjoint.
@@
Theorem VII.4.
Let $\mathbb{M}_{\infty},\mathbb{M}_{1},V,\mathcal{G}_{\infty},\mathcal{G}_{1}$ be defined as in Theorem VII.2. Suppose $\hat{\mu}_{\infty},\hat{\mu}_{1}$ are the nearest elements of $\mathbb{M}_{\infty},\mathbb{M}_{1}$ under the $V$-norm: $\hat{\mu}_{\infty},\hat{\mu}_{1}=\operatorname*{arg\,min}_{\mu_{\infty}\in\mathbb{M}_{\infty},\mu_{1}\in\mathbb{M}_{1}}\|\mu_{\infty}-\mu_{1}\|_{V}^{2}$. Then, the distributions $G_{\hat{\mu}_{\infty}},G_{\hat{\mu}_{1}}$ are the nearest elements of $\mathcal{G}_{\infty},\mathcal{G}_{1}$ under the Fisher distance: $\mathbb{D}_{F}(G_{\hat{\mu}_{1}}\|G_{\hat{\mu}_{\infty}})=\min_{R_{\infty}\in\mathcal{G}_{\infty},R_{1}\in\mathcal{G}_{1}}\mathbb{D}_{F}(R_{1}\|R_{\infty})$.

### v3

@@ I Introduction
In contrast with the work in [17], we define the notion of least favorable distribution using Fisher divergence and provide a method to identify the least favorable distributions.
@@
    We propose a new robust score-based quickest change detection algorithm that can be applied to score-based models and unnormalized models, namely, statistical models whose density involves an unknown normalizing constant. We generalize the SCUSUM algorithm of [15] to propose a robust score-based variant which we refer to as RSCUSUM. As with the SCUSUM method, the role of Kullback-Leibler divergence in classical change detection is replaced with the Fisher divergence between the pre-and post-change distributions, and the role of log-likelihood ratios is replaced with differences between Hyvärinen scores ([16]). This algorithm is introduced in Section V.
@@
    As with the SCUSUM algorithm, the RSCUSUM algorithm is a score-based algorithm where cumulative scores do not enjoy a standard martingale characterization. This differentiates RSCUSUM and SCUSUM from the classic CUSUM method, which leverages the fact that the likelihood ratios form a martingale under the pre-change model [9, 18]. We generalize the analysis of the delay and false-alarm of the SCUSUM method to the robust case. This analysis is presented in Section VI.
@@ II Problem Formulation
$\text{ARL}(T)\overset{\Delta}{=}\mathbb{E}_{\infty}[T].$
@@
$\min_{T}\;\sup_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathcal{L}_{\texttt{WADD}}(T)\;\quad\text{subject to}\;\quad\inf_{P_{\infty}\in\mathcal{G}_{\infty}}\mathbb{E}_{\infty}[T]\geq\gamma,$ (2)
where $\gamma$ is a constraint on the ARL. The delay $\mathcal{L}_{\texttt{WADD}}$ in the above problem is a function of the true pre- and post-change laws $P_{\infty},P_{1}$ and should be designated as $\mathcal{L}_{\texttt{WADD}}^{P_{\infty},P_{1}}.$ We will, however, suppress this notation and refer to $\mathcal{L}_{\texttt{WADD}}^{P_{\infty},P_{1}}$ by $\mathcal{L}_{\texttt{WADD}}$. Similarly, we will suppress the dependence of $\mathbb{E}_{\infty}[T]$ on the pre-change law $P_{\infty}$. Thus, the goal in this problem is to find a stopping time $T$ to minimize the worst-case detection delay, subject to a constraint $\gamma$ on $\mathbb{E}_{\infty}[T]$. We are also interested in the minimax formulation:
$\min_{T}\;\sup_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathcal{L}_{\texttt{CADD}}(T)\;\quad\text{subject to}\;\quad\inf_{P_{\infty}\in\mathcal{G}_{\infty}}\mathbb{E}_{\infty}[T]\geq\gamma.$ (3)
@@ IV Quickest Change Detection in Unnormalized and Score-Based Models
whenever it can be well defined. We assume that the Hyvärinen scores is well defined for all $P\in\mathcal{G}_{\infty}\cup\mathcal{G}_{1}$ by Assumption IV.4. Here, $\|\cdot\|_{2}$ denotes the Euclidean norm, $\nabla_{x}$ and $\Delta_{x}=\sum_{i=1}^{d}\frac{\partial^{2}}{\partial x_{i}^{2}}$ respectively denote the gradient and the Laplacian operators acting on $X=(x_{1},\cdots,x_{d})^{\top}$.
@@
$\mathbb{D}_{\texttt{F}}(P\|Q)\overset{\Delta}{=}\mathbb{E}_{X\sim P}\left[\frac{1}{2}\left\|\nabla_{{x}}\log p(X)-\nabla_{{x}}\log q(X)\right\|_{2}^{2}\right],$ (8)
@@
where $\omega>0$ is a stopping threshold that is pre-selected to control false alarms, and $Y(n)$ can be computed recursively:
@@
We now impose the following conditions on the classes $\mathcal{G}_{\infty}$ and $\mathcal{G}_{1}$ in order to make claims regarding the SCUSUM Algorithm. All lemmas and theorems in the paper, unless otherwise stated, will be assumed to be valid when all these conditions are satisfied.
Assumption IV.3.
$\mathcal{G}_{\infty},\mathcal{G}_{1}$ are disjoint and are each convex.
Assumption IV.4.
Following regularity conditions introduced in [16] hold:
  1. 1.
    For all densities $g\in\mathcal{G}_{\infty}\cup\mathcal{G}_{1}$, $g$ is differentiable.
  2. 2.
    For all $g\in\mathcal{G}_{\infty}\cup\mathcal{G}_{1}$, $\mathbb{E}_{x\sim g}[\|\nabla_{x}\log g(x)\|^{2}_{2}]<\infty$.
  3. 3.
    For all $g\in\mathcal{G}_{\infty}\cup\mathcal{G}_{1}$ and for all $h\in\mathcal{G}_{\infty}\cup\mathcal{G}_{1}$, $g(x)\nabla_{x}\log h(x)\rightarrow 0$ as $\|x\|\rightarrow\infty$.
Assumption IV.5.
Note that the regularity conditions contained in Assumption IV.4 were assumed in [16] to ensure that the Hyvärinen score is well-defined. We also assume that all the moment-generating functions appearing in the paper are finite. The following theorem is established in [15] regarding the consistency and the performance of the SCUSUM algorithm.
Theorem IV.6 ([15]).
The following statements are true for the SCUSUM algorithm:
@@ V Robust Quickest Change Detection in Unnormalized and Score-Based Models
$(Q_{\infty},Q_{1})=\operatorname*{arg\,min}_{(R_{\infty},R_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{F}(R_{1}\|R_{\infty}).$ (14)
@@
The motivation for using (14) comes from the following analysis. Note that for any stopping rule $T$ such that $\inf_{P_{\infty}\in\mathcal{G}_{\infty}}\mathbb{E}_{\infty}[T]\geq\gamma$, we have as $\gamma\to\infty$,
$\min_{T}\;\sup_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathcal{L}_{\texttt{WADD}}(T)\geq\sup_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\min_{T}\;\mathcal{L}_{\texttt{WADD}}(T)\geq\frac{\log\gamma}{\inf_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{KL}(P_{1}\|P_{\infty})}(1+o(1)),$ (15)
where the $o(1)$ terms goes to zero as $\gamma\to\infty$. Thus,
$\frac{\log\gamma}{\inf_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{KL}(P_{1}\|P_{\infty})}$
provides a lower bound on the asymptotic performance of any algorithm. This lower bound can be achieved by a CUSUM-type algorithm $T^{*}$ if $\inf_{P_{\infty}\in\mathcal{G}_{\infty}}\mathbb{E}_{\infty}[T^{*}]\geq\gamma$ and the drift of this algorithm after the change is $\inf_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{KL}(P_{1}\|P_{\infty})$. Since we cannot calculate the KL divergence for unnormalized and score-based models, we use the Fisher divergence as a proxy measure. We note also that these two divergences coincide for certain Gaussian families [15].
@@ VI Consistency and Delay and False Alarm Analysis of the RSCUSUM Algorithm
In this section, we prove the consistency (the ability to detect the change with probability one) and provide delay and false alarm analysis of the RSCUSUM algorithm. In Section VI-A, we prove an important lemma that can be interpreted as a reverse triangle inequality for Fisher divergence. This lemma is then used in Section VI-B to prove the consistency of the RSCUSUM algorithm. In Section VI-C, we obtain a lower bound on the mean time to a false alarm, and in Section VI-D, we obtain an expression for the average detection delay of the RSCUSUM algorithm.
We first make another fundamental assumption:
@@
If $\mathcal{G}_{\infty}$ contains only one element ($\mathcal{G}_{\infty}=\{P_{\infty}\}$), then clearly the least favorable pre-change distribution $Q_{\infty}$ is equal to $P_{\infty}$ and $\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})=0$. Then, Assumption VI.1 follows from the disjointness of $\mathcal{G}_{\infty},\mathcal{G}_{1}$ given in Assumption IV.3 and the regularity conditions given in Assumption IV.4.
@@
$\mathcal{P}_{m}=\{P_{i},\;i=1,\dots,m\},\;m\in\mathbb{N}^{+}$
where each distribution $P_{i}$ has density $p_{i}$. Further suppose that $\mathcal{G}_{\infty}$ is defined to be the convex hull of this finite set:
@@ VI-A Reverse Triangle Inequality for Fisher Divergence
$=\int\frac{1}{2}\big{\|}\nabla\log\bigl{(}\xi q_{1}+(1-\xi)r_{1}\bigr{)}-\nabla\log q_{\infty}\big{\|}^{2}$
@@
$\mathcal{L}^{\prime}(\xi)=\int\frac{1}{2}\bigl{(}q_{1}-r_{1}\bigr{)}\big{\|}\nabla\log q_{\xi}-\nabla\log q_{\infty}\big{\|}^{2}dx+\int q_{\xi}\nabla\left(\frac{q_{1}-r_{1}}{q_{\xi}}\right)^{T}\bigl{(}\nabla\log q_{\xi}-\nabla\log q_{\infty}\bigr{)}dx.$
@@
$\mathcal{L}^{\prime}(1^{-})=\int\frac{1}{2}\bigl{(}q_{1}-r_{1}\bigr{)}\big{\|}\nabla\log q_{1}-\nabla\log q_{\infty}\big{\|}^{2}dx\quad+\int q_{1}\nabla\left(\frac{q_{1}-r_{1}}{q_{1}}\right)^{T}\bigl{(}\nabla\log q_{1}-\nabla\log q_{\infty}\bigr{)}dx$
$=\mathbb{D}_{\texttt{F}}\left(Q_{1}\|Q_{\infty}\right)-\int\underbrace{\frac{1}{2}r_{1}\big{\|}\nabla\log q_{1}-\nabla\log q_{\infty}\big{\|}^{2}}_{\text{term 1}}+\underbrace{q_{1}\nabla\left(\frac{q_{1}-r_{1}}{q_{1}}\right)^{T}\bigl{(}\nabla\log q_{1}-\nabla\log q_{\infty}\bigr{)}}_{\text{term 2}}dx.$ (20)
@@
$\frac{1}{2}r_{1}\big{\|}\nabla\log q_{1}-\nabla\log q_{\infty}\big{\|}^{2}$
$=\frac{1}{2}r_{1}\big{\|}\nabla\log q_{1}-\nabla\log r_{1}\big{\|}^{2}+\frac{1}{2}r_{1}\big{\|}\nabla\log r_{1}-\nabla\log q_{\infty}\big{\|}^{2}+\underbrace{r_{1}\bigl{(}\nabla\log q_{1}-\nabla\log r_{1}\bigr{)}^{T}\bigl{(}\nabla\log r_{1}-\nabla\log q_{\infty}\bigr{)}}_{\text{term 1(a)}}.$ (21)
@@
$q_{1}\nabla\left(\frac{q_{1}-r_{1}}{q_{1}}\right)^{T}\bigl{(}\nabla\log q_{1}-\nabla\log q_{\infty}\bigr{)}=r_{1}\bigl{(}\nabla\log q_{1}-\nabla\log r_{1}\bigr{)}^{T}\bigl{(}\nabla\log q_{1}-\nabla\log q_{\infty}\bigr{)}.$ (24)
@@ VI-C False Alarm Analysis of the RSCUSUM Algorithm
We emphasize that $\rho$ is a quantity that depends upon $P_{\infty},Q_{\infty}$, and $Q_{1}$ but which does not depend upon $\omega$. For any choice of $\mathcal{G}_{\infty},\mathcal{G}_{1},P_{\infty}$ that satisfies the assumptions of this paper, we can show the existence of a $\rho$ that is a solution to (26).
@@
Theorem VI.7 implies that the ARL increases at least exponentially as the stopping threshold $\omega$ increases. Since the bound is valid for all those $\rho$ values for which $\mathbb{E}_{\infty}[\exp(z_{\rho}(X_{1}))]\leq 1$, the bound is valid for the biggest such number, which is $\rho^{*}$. The existence of such a $\rho^{*}$ is guaranteed by Lemma VI.6. Thus, we have
$\mathbb{E}_{\infty}[T_{\texttt{RSCUSUM}}]\geq e^{\rho^{*}\omega}.$ (28)
As a result, by setting $\omega=\frac{\log\gamma}{\rho^{*}}$, we get
$\mathbb{E}_{\infty}[T_{\texttt{RSCUSUM}}]\geq\gamma.$ (29)
We note again that $\rho^{*}$ depends on $P_{\infty},Q_{\infty}$, and $Q_{1}$.
@@ VI-D Delay Analysis of the RSCUSUM Algorithm
Theorem VI.8 implies that the expected detection delay (EDD) increases at most linearly as the stop threshold $\omega$ increases for large values of $\omega$. Again, since the bound is valid for all those $\rho$ values for which $\mathbb{E}_{\infty}[\exp(z_{\rho}(X_{1}))]\leq 1$, the bound is valid for the biggest such number, which is $\rho^{*}$. This gives
$\mathcal{L}_{\texttt{WADD}}(T_{\texttt{RSCUSUM}})\sim\mathcal{L}_{\texttt{CADD}}(T_{\texttt{RSCUSUM}})\lesssim\frac{\log\gamma}{\rho^{*}\mathbb{D}_{F}(Q_{1}\|Q_{\infty})},\quad\text{as $\gamma\to\infty$.}$ (30)
@@ VII Identification of the Least Favorable Distributions
$\mathbb{E}_{P_{*}}\biggl{\{}\frac{1}{2}\|\nabla_{x}\log p_{*}(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{\}}=\min_{p\in\mathcal{A}_{m},\phi\in\mathcal{B}_{m}}\mathbb{E}_{P}\biggl{\{}\frac{1}{2}\|\phi(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{\}}.$ (33)
@@
$\mathbb{E}_{P_{*}}\biggl{\{}\frac{1}{2}\|\nabla_{x}\log p_{*}(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{\}}=\min_{p\in\mathcal{A}_{m}}\mathbb{E}_{P}\biggl{\{}\frac{1}{2}\|\nabla_{x}\log p(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{\}}.$
@@
$\mathbb{E}_{P}\biggl{\{}\frac{1}{2}\|\nabla_{x}\log p(X)-\nabla_{x}\log r(X)\|_{2}^{2}\biggr{\}}=\mathbb{E}_{P}\biggl{\{}\frac{1}{2}\biggl{\|}\frac{\nabla_{x}p(X)}{p(X)}-\nabla_{x}\log r(X)\biggr{\|}_{2}^{2}\biggr{\}}$
$=\mathbb{E}_{P}\biggl{\{}\frac{1}{2}\biggl{\|}\frac{\sum_{i=1}^{m}w_{i}\nabla_{x}p_{i}(X)}{\sum_{i=1}^{m}w_{i}p_{i}(X)}-\nabla_{x}\log r(X)\biggr{\|}_{2}^{2}\biggr{\}}=\mathbb{E}_{P}\biggl{\{}\frac{1}{2}\biggl{\|}\sum_{i=1}^{m}u_{i}(X)\nabla_{x}\log p_{i}(X)-\nabla_{x}\log r(X)\biggr{\|}_{2}^{2}\biggr{\}},$
@@
Next, we provide a method to find the LFD in a class of Gaussian mixture models. We begin by defining a $V$-norm:
Definition VII.2 ($V$-norm).
The $V$-norm of a vector $\theta$ is given as
$\|\theta\|_{V}\overset{\Delta}{=}\sqrt{\theta^{T}V^{-2}\theta}$ (34)
Theorem VII.3.
Let $\mathbb{M}_{\infty},\mathbb{M}_{1}\subset\mathbb{R}^{d}$ be disjoint, convex, and compact sets in $d-$dimensional Euclidean space. Fix a symmetric, positive-definite matrix $V\in\mathbb{R}^{d\times d}$. For $\theta\in\mathbb{M}_{\infty}\cup\mathbb{M}_{1}$, let $G_{\theta}$ be the Gaussian distribution with density $g_{\theta}$ parameterized by covariance matrix $V$ and mean $\theta:G_{\theta}=\mathcal{N}(\theta,V)$. Let the distribution class $\mathcal{G}_{\infty}$ be the convex hull of all Gaussian $\{G_{\theta}:\theta\in\mathbb{M}_{\infty}\}$ and let $\mathcal{G}_{1}$ be the corresponding convex hull of $\{G_{\theta}:\theta\in\mathbb{M}_{1}\}$. Then, $\mathcal{G}_{\infty},\mathcal{G}_{1}$ are disjoint.
@@
Theorem VII.5.
Let $\mathbb{M}_{\infty},\mathbb{M}_{1},V,\mathcal{G}_{\infty},\mathcal{G}_{1}$ be defined as in Theorem VII.3. Suppose $\hat{\mu}_{\infty},\hat{\mu}_{1}$ are the nearest elements of $\mathbb{M}_{\infty},\mathbb{M}_{1}$ under the $V$-norm: $(\hat{\mu}_{\infty},\hat{\mu}_{1})=\operatorname*{arg\,min}_{(\mu_{\infty},\mu_{1})\in\mathbb{M}_{\infty}\times\mathbb{M}_{1}}\|\mu_{\infty}-\mu_{1}\|_{V}^{2}$. Then, the distributions $G_{\hat{\mu}_{\infty}},G_{\hat{\mu}_{1}}$ are the nearest elements of $\mathcal{G}_{\infty},\mathcal{G}_{1}$ under the Fisher distance: $\mathbb{D}_{F}(G_{\hat{\mu}_{1}}\|G_{\hat{\mu}_{\infty}})=\min_{(R_{\infty},R_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{F}(R_{1}\|R_{\infty})$.